OpenStudy (anonymous):

What is the probability of randomly selecting either a club or a non-face card from a standard deck of cards?

OpenStudy (anonymous):

was thinking: 13/ 52 + 40/52 but the answer is 43/52

OpenStudy (anonymous):

9 are face cards that are not clubs so 43 remain.

OpenStudy (anonymous):

that means \[\frac{43}{52}\]

OpenStudy (anonymous):

wait does ace count as face card?

OpenStudy (anonymous):

no faces on aces

OpenStudy (anonymous):

so its just jack, queen, king?

OpenStudy (anonymous):

yes those

OpenStudy (anonymous):

how did u get 9? 13-3 = 10?...

OpenStudy (anonymous):

club or non face card. so only ones that are out are the jack queen and king of hearts, spades and clubs

OpenStudy (anonymous):

so only 9 cards are not included. this leaves 52-9=43 that are either clubs or not face cards

OpenStudy (anonymous):

oh wow thx

OpenStudy (anonymous):

yw